OpenStudy (anonymous):

Natalie makes $2,000 per month. She spends $100 on credit card payments and $250 on an auto loan. What is her debt-to-income ratio?

OpenStudy (bahrom7893):

2000/350 ? is that it?

OpenStudy (anonymous):

The annual percentage rate 13.5

OpenStudy (anonymous):

= 35/ 200 = 7/40